# DaGRPO: Rectifying Gradient Conflict in Reasoning via Distinctiveness-Aware Group Relative Policy Optimization

### v1

@@ 2 Related Work
Optimization and Analysis of GRPO. As the core algorithm of DeepSeek-R1 (guo2025deepseek), Group Relative Policy Optimization (GRPO) replaces the traditional value function with group-wise relative advantage estimation, demonstrating a superior performance-compute ratio in mathematical and code reasoning tasks. However, its widespread adoption has exposed inherent limitations, catalyzing a surge of targeted research. Addressing training instability, DAPO (yu2025dapo) identifies the risk of entropy collapse and proposes clip-higher and dynamic sampling strategies, GSPO (zheng2025group) improves the stability of MoE model training by elevating the optimization granularity from the token level to the sequence level, while GTPO (simoni2025gtpo) highlight the phenomenon of gradient conflict and attempt to mitigate its impact on critical reasoning steps through token-level gradient correction. Concurrently, wu2025takes reformulate the GRPO objective from a contrastive learning perspective, revealing its intrinsic mathematical connection with DPO. To further tackle the sample inefficiency of exclusive on-policy sampling, recent works have explored hybrid paradigms incorporating off-policy data. For instance, CHORD (zhang2025policy) utilizes a global coefficient to dynamically balance the gradient weights between SFT data (off-policy) and GRPO samples (on-policy), while LUFFY (yan2025learning) integrates offline trajectories directly into online groups. Different from these approaches, we identify the lack of sample distinctiveness as the root cause of optimization instability and propose a distinctiveness-aware mechanism that dynamically filters homogeneous conflicts and introduces high-quality contrastive references.
@@ 5 Method
This hybrid objective effectively synergizes the exploration capability of on-policy RL with the guidance of off-policy demonstrations. By explicitly rectifying the gradient direction through distinctiveness masking and anchoring the optimization trajectory via high-quality demonstrations, DaGRPO mitigates the variance induced by token-level ambiguity, leading to more stable convergence in complex mathematical domains.
@@ 6 Experimental Setups
  • OOD Generalization (3 Datasets): To verify that our method improves general reasoning beyond math without catastrophic forgetting, we evaluate on ARC–c (clark2018think) (grade–school science), GPQA–Diamond (rein2024gpqa) (expert–level biology, physics, and chemistry), and MMLU–Pro (wang2024mmlu) (interdisciplinary complex reasoning across 57+ subjects).
@@ 7.1 Main Results
Robust Out–of–Distribution Generalization. Beyond mathematical reasoning, DaGRPO demonstrates exceptional robustness on OOD tasks. Both DaGRPO variants achieve superior generalization performance compared to all baselines. Notably, DaGRPO (w/o Off-policy) attains the highest average OOD score of 59.9%, significantly outperforming GRPO (57.2%). This suggests that our distinctiveness–aware update not only enhances reasoning but also prevents the “alignment tax” or overfitting often associated with aggressive RL fine–tuning, preserving the model’s broad scientific capabilities.
@@ 8 Conclusion
Despite its promising performance, DaGRPO has certain limitations. First, the gradient rectification mechanism relies on an LLM-as-a-Judge for fine-grained scoring, which introduces additional computational overhead and latency during the training phase. While this does not affect inference, future work could explore distilling lightweight reward models to replace the LLM judge for more efficient training. Second, the off-policy augmentation component depends on high-quality external expert data (e.g., from DeepSeek-R1) acting as anchors. In scenarios where strong prior models are unavailable, generating high-distinctiveness sample pairs solely through self-bootstrapping remains a challenging open problem. We leave these explorations for future work to further advance efficient and autonomous reasoning alignment.

### v2

@@ 2 Related Work
Optimization and Analysis of GRPO. As the core algorithm of DeepSeek-R1 (guo2025deepseek), Group Relative Policy Optimization (GRPO) replaces the traditional value function with group-wise relative advantage estimation, demonstrating a superior performance-compute ratio in mathematical and code reasoning tasks. However, its widespread adoption has exposed inherent limitations, catalyzing a surge of targeted research. Addressing training instability, DAPO (yu2025dapo) identifies the risk of entropy collapse and proposes clip-higher and dynamic sampling strategies, GSPO (zheng2025group) improves the stability of MoE model training by elevating the optimization granularity from the token level to the sequence level. GTPO (simoni2025gtpo) attempts to mitigate gradient conflict by selectively masking the gradient contributions of conflicting tokens. However, this strategy typically only identifies continuous conflicting tokens at the sequence boundaries (i.e., beginning and end), effectively ignoring the conflicts triggered by insufficient logical distinctiveness deep within the reasoning chains. Concurrently, wu2025takes reformulate the GRPO objective from a contrastive learning perspective, revealing its intrinsic mathematical connection with DPO. To further tackle the sample inefficiency of exclusive on-policy sampling, recent works have explored hybrid paradigms incorporating off-policy data. For instance, CHORD (zhang2025policy) utilizes a global coefficient to dynamically balance the gradient weights between SFT data (off-policy) and GRPO samples (on-policy), while LUFFY (yan2025learning) integrates offline trajectories directly into online groups. Different from these approaches, we identify the lack of sample distinctiveness as the root cause of optimization instability and propose a distinctiveness-aware mechanism that dynamically filters homogeneous conflicts and introduces high-quality contrastive references.
@@ 5 Method
Crucially, the distinctiveness-aware masking described in Eq. (13) applies to the mixed group $G_{\text{mix}}$. This provides a safety mechanism against noisy or low-quality off-policy data. If an injected anchor $o_{\text{off}}$ is factually incorrect (low score) or only marginally better than the current policy’s best response (score gap $\delta$), the mask $\lambda$ will automatically zero out its contribution. Consequently, DaGRPO is robust to the quality variance of external demonstrators, ensuring that the policy only learns from anchors that provide a strictly superior and unambiguous reasoning signal.
@@ 6 Experimental Setups
  • OOD Generalization (3 Datasets): To verify that our method improves general reasoning beyond math without catastrophic forgetting, we evaluate on ARC–c (clark2018think), GPQA–Diamond (rein2024gpqa) (expert–level biology, physics, and chemistry), and MMLU–Pro (wang2024mmlu) (interdisciplinary complex reasoning across 57+ subjects).
@@ 7.1 Main Results
Robust Out–of–Distribution Generalization. A key observation is that DaGRPO significantly outperforms baselines on OOD tasks (Avg 59.9% vs. GRPO 57.2%). We attribute this to the preservation of policy entropy shown in Figure 2 (Right). Standard GRPO tends to collapse into narrow, repetitive solution templates (low entropy) to maximize rewards on in-distribution math problems, which harms generalization. By masking out ambiguous samples that encourage “safe” but mediocre responses, DaGRPO maintains a higher exploration entropy, preventing overfitting to specific prompt-response patterns and preserving the model’s broad reasoning capabilities for unseen domains like biology (GPQA) and science (ARC-c).
@@ 8 Conclusion
Despite its promising performance, DaGRPO has certain limitations. First, the gradient rectification mechanism relies on an LLM-as-a-Judge for fine-grained scoring, which introduces additional computational overhead and latency during the training phase. However, this is not a permanent bottleneck; future work can distill the judge’s scoring capability into a lightweight, regression-based Reward Model (RM) to replace the online API calls, significantly reducing training latency while maintaining distinctiveness sensitivity. Second, the off-policy augmentation component depends on high-quality external expert data (e.g., from DeepSeek-R1) acting as anchors. In scenarios where strong prior models are unavailable, generating high-distinctiveness sample pairs solely through self-bootstrapping remains a challenging open problem. We leave these explorations for future work to further advance efficient and autonomous reasoning alignment.